Heptane is the zero of the octane rating of motor fuel. It burns according to the following:
C7H16 + 11O2 à 7CO2 + 8H2O
Estimate the enthalpy change in the reaction by using bond dissociation energies, and classify the reaction as exothermic or endothermic overall. [Average BDEs in kcal/mol: C-H, 99; O=O, 119; C=O, 191; H-O 112]. (Hint: C7H16 has 16 C-H bonds)
Answer #1

+ 1 1 CHIO + 11 O2 → 7602 +8H2O Bond broken = 161C-n) + 6(0-c) + 11 (0:0) = 16 x 99 + 6 x 83 + 11 x 119 = 3391 kcal/mol Bond